References

[1] E.D.I. Thor & E. di Thor, editors (1987): Title B1: A Book with editors, Series 443. Publisher, Address. Note.

[2] E.D.I. Thor & E. di Thor, editors (1987): Title B2: An Inbook with editors, Series 443, type (of chapter) II, pages 1–999. Publisher, Address. Note.

[3] E.D.I. Thor & E. di Thor, editors (1987): Title B3: A Proceedings, Series 443, Publisher, Address. Note.

[4] Title B3: A Proceedings without editor or organization, Series 443, Publisher, Address, Month 1987. Note.

[5] A. Ut, H. Or & Co Author (1987): Title a: An article. Jounal 443(21), pp. 1–999. Note.

[6] A. Ut, H. Or & Co Author (1987): Title B1: A Book with authors, Series 443. Publisher, Address. Note.

[7] A. Ut, H. Or & Co Author (1987): Title B2: An Inbook with authors, Series 443, type (of chapter) II, pages 1–999. Publisher, Address. Note.

[8] A. Ut, H. Or & Co Author (1987): Title c1: An inproceedings. In E.D.I. Thor & E. di Thor, editors: Booktitle, Series 443, Publisher, Address, pp. 1–999. Note.

[9] A. Ut, H. Or & Co Author (1987): Title c2: An incollection. In E.D.I. Thor & E. di Thor, editors: Booktitle, Series 443, type II, Publisher, Address, pp. 1–999. Note.

[10] A. Ut, H. Or & Co Author (1987): Title r: A techreport. Type (of report) 21, Institution, Address. Note.

[11] A. Ut, H. Or & Co Author (1987): Title r1: A techreport. Technical Report 21, Institution, Address. Note.

[12] A. Ut, H. Or & Co Author (1987): Title R2: A PhDthesis. PhD thesis, School, Address. Note.

[13] A. Ut, H. Or & Co Author (1987): Title r3: A masterthesis. Master’s thesis, School, Address. Note.

[14] A. Ut, H. Or & Co Author (1987): Title R4: A Manual. Organization, Address. Note.

[15] A. Ut, H. Or & Co Author (1987): Title u: A booklet. Howpublished, Address. Note.

[16] A. Ut, H. Or & Co Author (1987): Title u: A misc. Howpublished. Note.

[17] A. Ut, H. Or & Co Author (1987): Title u: An unpublished. Note.